# MVP-ORAM: a Wait-free Concurrent ORAM for Confidential BFT Storage

### v1

@@ F-E3 Experiment (E3)
[Comparison] [5 human-minutes + 1 compute-hour]: This experiment shows that MVP-ORAM outperforms QuORAM both in throughput and latency with up to $50$ clients. However, due to the resources required to run experiments, we only show results for up to $15$ clients.
@@
[Execution] Navigate to MVP-ORAM-Artifact/QuORAM in builder terminal, and then follow the steps described in the experimental workflow to execute the experiment using the above values.
@@
gnuplot -e "O=’<output.path>’; L=’17’; Z=’4’; B=’8’; A=’1.0’; c_max=’5’" plotScripts/mvp_oram_vs_quoram_plot.gp
The correct execution of this command will create the <output.path>/output/plots/comparison.pdf file containing a figure comparing MVP-ORAM with QuORAM. Due to the scaled-down experiment, this figure only shows the overall trend of Table III, confirming (C4).

### v2

@@ F-E3 Experiment (E3)
[Comparison with QuORAM] [5 human-minutes + 1 compute-hour]: This experiment shows that MVP-ORAM outperforms QuORAM both in throughput and latency. However, due to the resources required to run the experiments, we only show results for up to $15$ clients.
@@
[Execution] Consider the project QuORAM and follow the steps described in the experimental workflow to execute the experiment using the above values.
@@
gnuplot -e "O=’<output.path>’; L=’16’; Z=’4’; B=’8’; A=’1.0’; c_max=’5’" plotScripts/mvp_oram_vs_quoram_plot.gp
The correct execution of this command will create the <output.path>/output/plots/quoram.pdf file containing a figure comparing MVP-ORAM with QuORAM. Due to the scaled-down experiment, this figure only shows the overall trend of Table IV, confirming (C4).
F-E4 Experiment (E4)
[Comparison with COBRA] [5 human-minutes + 0.3 compute-hour]: This experiment shows that MVP-ORAM’s performance is an order of magnitude lower than COBRA with up to $50$ clients. However, due to the resources required to run the experiments, we present results only for up to $15$ clients.
[Preparation] Use the following parameters for this experiment:
  • global.worker.machines=9
  • controller.benchmark.strategy=oram. benchmark.cobra.ThroughputLatencyBench markStrategy
  • fault_thresholds=1 2
  • clients_per_round=1 5 10 15
  • cobra.request.private_data_size=8
  • measurement_duration=60
[Execution] Consider the project MVP-ORAM and follow the steps described in the experimental workflow to execute the experiment using the above values.
[Results] Execute the following command in the builder terminal:
gnuplot -e "O=’<output.path>’; L=’16’; Z=’4’; B=’8’; A=’1.0’; c_max=’5’" plotScripts/mvp_oram_vs_cobra_plot.gp
The correct execution of this command will create <output.path>/output/plots/cobra.pdf file containing a figure comparing MVP-ORAM with COBRA. Due to the scaled-down experiment, this figure only shows the overall trend of Table IV, confirming (C5).
F-E5 Experiment (E5)
[Impact of $\alpha$ and $L$ on security] [10 human-minutes + 0.01 compute-hour]: This numerical experiment demonstrates the security improvements of MVP-ORAM when the access distribution is skewed and the tree size increases.
[Preparation] This experiment requires executing the steps described in the experimental workflow twice, with different values of tree_heights and access_thresholds. In both executions, use the following parameters:
  • global.worker.machines=0
  • controller.benchmark.strategy=oram. benchmark.tvd.TVDHeightBenchmarkStrategy
  • clients_per_round=1 5 10 15 20 25 30 35 40 45 50
First execution: Set tree_heights=17 and access_thresholds=(0.90436, 0.8) (1.0945, 0.9) (1.2353, 0.95) (1.537, 0.99).
Second execution: Set tree_heights=25 and access_thresholds=(0.87683, 0.8) (1.0251, 0.9) (1.1237, 0.95) (1.3101, 0.99).
[Execution] Consider the project MVP-ORAM and follow the steps, except the last step, described in the experimental workflow to execute the experiment using the above values.
gnuplot -e "O=’<output.path>’; L1=’17’; L2=’25’" plotScripts/tvd_heights_plot.gp
The correct execution of this command will create <output.path>/output/plots/tvd_heights.pdf file containing the figure, confirming (C6).
F-E6 Experiment (E6)
[Impact of $\sigma$ on security] [5 human-minutes + 0.01 compute-hour]: This numerical experiment demonstrates the security improvements of Strong MVP-ORAM as we increase the number of dummy accesses $\sigma$.
[Preparation] Use the following parameters for this experiment:
  • global.worker.machines=0
  • controller.benchmark.strategy=oram. benchmark.tvd.TVDSigmaBenchmarkStrategy
  • clients_per_round=1 5 10 15 20 25 30 35 40 45 50
  • tree_heights=17
  • access_thresholds=(0.0000001, 0.99) (1.537, 0.99)
  • sigmas=0 10 20 30 40 50
[Execution] Consider the project MVP-ORAM and follow the steps, except the last step, described in the experimental workflow to execute the experiment using the above values.
gnuplot -e "O=’<output.path>’; L=’17’; T=’0.99’" plotScripts/tvd_sigmas_plot.gp
The correct execution of this command will create <output.path>/output/plots/tvd_sigmas.pdf file containing the figure, confirming (C7).